# LBMamba: Locally Bi-directional Mamba

### v1

@@ 4 Experiments
In this section, we present a series of experiments to evaluate the performance of LBVim and compare it mainly to DeiT (Touvron et al., 2021) and Vim (Zhu et al., 2024) across various visual tasks. We also apply LBMamba to SOTA MIL method MambaMIL and SRMambaMIL to evaluate its performance on WSI datasets (Yang et al., 2024b). Following Mamba(Gu & Dao, 2023), we set the number of elements each thread thread process ($M$) based on the sequence length $L$: when $L>256$ (images larger than $256\times 256$), a thread processes $M=16$ elements; for $128<L\leq 256$ (images between $256\times 256$ and $256\times 128$), it processes $M=8$ elements; and when $L\leq 128$, the workload is reduced to $M=4$. All natural image models are trained on 2/4 Nvidia A100/H100 GPUs. All throughput, GPU memory consumption and pathology models are run and evaluated on a Nvidia Quadro RTX 8000 GPU. Other implementation details are in Appendix.
@@ 4.1 Natural image classification
We first evaluate LBVim on the ImageNet-1K dataset (Deng et al., 2009), which contains 1.28M training images and 50K validation images from 1,000 categories. All models are trained on the training set, and top-1 accuracy on the validation set is reported. For fair comparisons, we follow the training setting in (Zhu et al., 2024). To be specific, we train our models for 300 epochs using a batch size of 1,024 for tiny model and a batch size of 512 otherwise. We use the AdamW optimizer (Loshchilov & Hutter, 2017) with a momentum of 0.9, a cosine annealing learning with an initial value of $1\times 10^{-3}$, a 5-epoch warmup period and a weight decay of 0.05. For data augmentation, we apply standard techniques: random cropping, horizontal flipping, label-smoothing regularization, mixup, and random erasing.
@@
Compared with Vim, LBVim has better speed-accuracy trade-off. We introduce two intermediate model configurations, LBVim-300 and LBVim-528. Where 300 and 528 are the dimension of features. As shown in table 1, their throughput closely match those of Vim-T and Vim-S, but surpassing the corresponding baselines by 1.6% and 0.8% in accuracy, respectively. As illustrated in figure 4, the accuracy–throughput curve of our models consistently lies in the upper-right quadrant relative to Vim, highlighting a more favorable trade-off across a wide operating range. These findings also show that instead of conducting an additional backward scan, it is better to spend the running time on larger model dimensions.
@@ 4.2 FLOPs and speed analysis
4.3 Downstream tasks on natural images
We evaluate the performance of LBVim on downstream tasks, including semantic segmentation on ADE20K dataset (Zhou et al., 2019), and object detection and instance segmentation on the COCO 2017 dataset (Lin et al., 2014). The training framework is based on the MMSegmenation (Contributors, 2020) and MMDetection (Chen et al., 2019) libraries, following (Zhu et al., 2024) in utilizing UperNet (Xiao et al., 2018) and Cascade Mask R-CNN (Cai & Vasconcelos, 2019) as the segmentation and detection networks, respectively. For a fair comparison, we add Linear layers to LBVim-300 and LBVim-528, adjusting their output dimension to be the same as Vim-Ti and Vim-S, respectively. We find that for these two downsteam tasks, which operates on dense features and therefore does not rely on a dedicated class token, the performance gap between LBVim and the Vim baseline narrows markedly.
Table 3: Left: The performance of our LBVim on the ADE20K semantic segmentation dataset. FLOPs and and throughput (T.P.) are measured with an input size of 512 × 2048. Right: The performance of LBVim on the COCO detection dataset (image are of size 1024$\times$1024). APb and APm denote the average precision for bonding boxes and masks, respectively. T.P. denotes average throughput.
@@ 4.4 Ablation study
We also conduct an ablation on the output aggregation scheme: global average pooling (GAP) versus multi-head attention pooling (MAP), across four capacity levels of LBVimin table 4 right. For the tiny, 300 and small variants, GAP delivers a 0.7%, 0.4% and 0.2% higher top-1 accuracy, respectively. It is clear that when the model scales up in dimension, the advantage of GAP narrows. When the model scales to LBVim-528, the MAP gains 1.1% at a negligible 0.7 % slowdown over GAP, indicating that the expressive benefit of learned pooling emerges only when sufficient parameters are available. This size-dependent trend aligns with prior observations that simple pooling is preferable for some compact models (Pan et al., 2021), whereas attention-based aggregation becomes advantageous in larger models (Dosovitskiy et al., 2020).
@@
$Acc$ $F1$ $AUC$ $Acc$ $F1$ $AUC$ $Acc$ $F1$ $AUC$
@@
4.5 WSI classification
To verify LBMamba beyond natural images, we embed it into the SOTA Multiple-Instance Learning (MIL) framework, MambaMIL and SRMambaMIL (Yang et al., 2024b), and name them LBMambaMIL and SRLBMambaMIL, respectively. We evaluate them on 3 public available Whole Slide Image datasets, PANDA (prostate grade assessment) (Bulten et al., 2022), TCGA-NSCLC (adenocarcinoma vs. squamous lung cancer) and TCGA-BRCA (breast invasive carcinoma sub-typing) (tcg, ). Dataset and training details are listed in Appendix D and E. As shown in table 5, with an additional locally backward scan, LBMambaMIL generally performans better than MambaMIL, achieves up to 3.06% higher accuracy, up to 3.39% higher F1 and up to 1.67% higher AUC. Similar improvement is also observed on SRMamba, SRLBMambaMIL achieves achieves up to 2.98% higher accuracy, up to 2.90% higher F1 and up to 1.48% higher AUC, demonstrating that the LBMamba is able to improve the performance of unidirectional scans. It also shows that our method generalize well on Giga-pixel pathology images.
@@ 4.5 WSI classification
The Effective Receptive Field (ERF) (Luo et al., 2016) refers to the region in the input space that contributes to the activation of a specific output unit. We conduct a comparative analysis of the central pixel’s ERF on DeiT, Vim and LBVim at tiny and small scale, both before and after training. The results presented in figure 5 illustrate that DeiT shows global ERFs but it suffers from the quadratic complexity of self attention. Vim shows global ERFs and LBVim also shows global ERFs, proving that LBMamba with sequence reversing operation does not have a side effect on ERF.
Figure 5: Comparison of Effective Receptive Fields (ERF) (Luo et al., 2016) on DeiT, Vim and LBVim at tiny and small scale. Pixels with higher intensity indicate larger responses related to the central pixel.
@@ 5 Conclusion
Limitation. LBMamba is less effective when a dedicated class token is appended to the sequence. We systematically evaluated four common variants: head class token (prepended), middle class token (inserted at the sequence midpoint), and double class token (both prepended and appended) strategy, but none of them achieves better performance than a simple global average pooling (see Appendix C). A plausible explanation is that the local backward scan treats the class token as an ordinary feature vector whose receptive field is confined to its local window; this reinforces local patterns while diluting the holistic summary that the token is meant to capture. Unifying local bi-directionality with an effective global summarization mechanism therefore remains an open research question, and we leave a deeper investigation of this to future work.
@@ Appendix A LBVim block algorithm
1:  /* normalize the input sequence $\mathbf{T}_{l-1}^{\prime}$/
2:  $\mathbf{T}_{l-1}^{\prime}$ : $(\mathtt{B},\mathtt{L},\mathtt{D})$leftarrow $\mathbf{Norm}(\mathbf{T}_{l-1})$
3:  $\mathbf{x}$ : $(\mathtt{B},\mathtt{L},\mathtt{E})$leftarrow $\mathbf{Linear}^{\mathbf{x}}(\mathbf{T}_{l-1}^{\prime})$
4:  $\mathbf{z}$ : $(\mathtt{B},\mathtt{L},\mathtt{E})$leftarrow $\mathbf{Linear}^{\mathbf{z}}(\mathbf{T}_{l-1}^{\prime})$
@@
6:  $\mathbf{x}^{\prime}$ : $(\mathtt{B},\mathtt{L},\mathtt{E})$leftarrow $\mathbf{SiLU}(\mathbf{Conv1d}_{o}(\mathbf{x}))$
7:  $\mathbf{B}$ : $(\mathtt{B},\mathtt{L},\mathtt{N})$leftarrow $\mathbf{Linear}^{\mathbf{B}}(\mathbf{x}^{\prime}_{o})$
8:  $\mathbf{C}$ : $(\mathtt{B},\mathtt{L},\mathtt{N})$leftarrow $\mathbf{Linear}^{\mathbf{C}}_{o}(\mathbf{x}^{\prime})$
9:  /* softplus ensures positive $\mathbf{\Delta}_{o}$/
10:  $\mathbf{\Delta}$ : $(\mathtt{B},\mathtt{L},\mathtt{E})$leftarrow $\log(1+\exp(\mathbf{Linear}^{\mathbf{\Delta}}(\mathbf{x}^{\prime})+\mathbf{Parameter}^{\mathbf{\Delta}}))$
11:  /* shape of $\mathbf{Parameter}^{\mathbf{A}}$ is $(\mathtt{E},\mathtt{N})$/
12:  $\overline{\mathbf{A}}$ : $(\mathtt{B},\mathtt{L},\mathtt{E},\mathtt{N})$leftarrow $\mathbf{\Delta}\bigotimes\mathbf{Parameter}^{\mathbf{A}}$
13:  $\overline{\mathbf{B}}$ : $(\mathtt{B},\mathtt{L},\mathtt{E},\mathtt{N})$leftarrow $\mathbf{\Delta}\bigotimes\mathbf{B}$
14:  /* initialization with 00 */
15:  $h^{f}$ : $(\mathtt{B},\mathtt{E},\mathtt{N})$leftarrow zeros $(\mathtt{B},\mathtt{E},\mathtt{N})$
16:  $h^{b}$ : $(\mathtt{B},\mathtt{E},\mathtt{N})$leftarrow zeros $(\mathtt{B},\mathtt{E},\mathtt{N})$
17:  $\mathbf{y}$ : $(\mathtt{B},\mathtt{L},\mathtt{E})$leftarrow zeros $(\mathtt{B},\mathtt{L},\mathtt{E})$
@@
32:  /* get gated $\mathbf{y}$/
33:  $\mathbf{y}^{\prime}$ : $(\mathtt{B},\mathtt{L},\mathtt{E})$leftarrow $\mathbf{y}\bigodot\mathbf{SiLU}(\mathbf{z})$
@@
35:  $\mathbf{T}_{l}$ : $(\mathtt{B},\mathtt{L},\mathtt{D})$leftarrow $\mathbf{Linear}^{\mathbf{T}}(\mathbf{y}^{\prime})+\mathbf{T}_{l-1}$
36:  $\mathbf{T}_{l}^{rev}$ : $(\mathtt{B},\mathtt{L},\mathtt{D})$leftarrow $\mathbf{Reverse}(\mathbf{T}_{l})$
@@ Appendix C Ablation on the class token
Table 7 shows several common strategies (Zhu et al., 2024) for incorporating a class token into LBVim-Ti on the ImageNet-100 dataset. Consistent with our observations on ImageNet-1K (section 4.1), LBVim-Ti under performs Vim-Ti by 0.42% as it can not process the class token. Introducing a head class token, which is prepended to the patch sequence (ViT style (Dosovitskiy et al., 2020)), reduces accuracy by 3.22%. Adding a second tail token (double class token) offers a recovery (0.58%) but remains 2.64% lower than GAP. Placing the token in the middle (the strategy used by Zhu et al. (2024)) of the sequence partially alleviates the degradation, yet it still under performs the no-token design. We believe this is because the class token is treated as an ordinary feature vector, its ability to aggregate global context is diluted, leading to weaker final representations. In contrast, GAP aggregates information from all positions without introducing additional parameters or disrupting the scan pattern, making it a more compatible summarization mechanism for LBMamba-based models.
@@
Appendix D Implementation details of WSI experiments

### v2

@@ 4 Experiments
In this section, we present a series of experiments to evaluate the performance of LBVim and compare it mainly to DeiT (touvron2021training_deit) and Vim (vim) across various visual tasks. We also apply LBMamba to SOTA MIL method MambaMIL and SRMambaMIL to evaluate its performance on WSI datasets (yang2024mambamil). Following Mamba(gu2023mamba), we set the number of elements each thread thread process ($M$) based on the sequence length $L$: when $L>256$ (images larger than $256\times 256$), a thread processes $M=16$ elements; for $128<L\leq 256$ (images between $256\times 256$ and $256\times 128$), it processes $M=8$ elements; and when $L\leq 128$, the workload is reduced to $M=4$. All natural image models are trained on 4/2 Nvidia A100/H100 GPUs. All throughput, GPU memory consumption and pathology models are run on a Nvidia Quadro RTX 8000 GPU. Other implementation details are in Appendix.
@@ 4.1 Natural image classification
We first evaluate LBVim on the ImageNet-1K dataset (deng2009imagenet), which contains 1.28M training images and 50K validation images from 1,000 categories. All models are trained on the training set, and top-1 accuracy on the validation set is reported. For fair comparisons, we follow the training setting in (vim). To be specific, we train our models for 300 epochs using a batch size of 1,024 for tiny model and a batch size of 512 otherwise. We use the AdamW optimizer (loshchilov2017decoupled_adamw) with a momentum of 0.9, a cosine annealing learning with an initial value of $1\times 10^{-3}$, a 5-epoch warmup period and a weight decay of 0.05. For data augmentation, we apply standard techniques: random cropping, horizontal flipping, label-smoothing regularization, mixup, and random erasing.
@@
Compared with Vim, LBVim has better speed-accuracy trade-off. We introduce two intermediate model configurations, LBVim-300 and LBVim-528. Where 300 and 528 are the dimension of features. As shown in table 1, their throughput closely match those of Vim-T and Vim-S, but surpassing the corresponding baselines by 1.6% and 0.8% in accuracy, respectively. As illustrated in figure 4, the accuracy–throughput curve of our models consistently lies in the upper-right quadrant relative to Vim, highlighting a more favorable trade-off across a wide operating range. These findings also show that our method enables a novel way to develop Mamba-based methods by scaling the model for higher accuracy, without slowing down the speed.
@@ 4.2 FLOPs and speed analysis
Table 3: Left: The performance of our LBVim on the ADE20K semantic segmentation dataset. FLOPs and and throughput (T.P.) are measured with an input size of 512 × 2048. Right: The performance of LBVim on the COCO detection dataset (image are of size $\times1024$). APb and APm denote the average precision for bonding boxes and masks, respectively. T.P. denotes average throughput.
@@
4.3 Downstream tasks on natural images
We evaluate the performance of LBVim on downstream tasks, including semantic segmentation on ADE20K dataset (zhou2019semantic_ade20k), and object detection and instance segmentation on the COCO 2017 dataset (lin2014microsoft_coco). The training framework is based on the MMSegmenation (mmseg2020) and MMDetection (mmdetection) libraries, following (vim) in utilizing UperNet (xiao2018unified_upernet) and Cascade Mask R-CNN (cai2019cascade_rcnn) as the segmentation and detection networks, respectively. For a fair comparison, we add Linear layers to LBVim-300 and LBVim-528, adjusting their output dimension to be the same as Vim-Ti and Vim-S, respectively. We find that for these two downsteam tasks, which operates on dense features and therefore does not rely on a dedicated class token, the performance gap between LBVim and the Vim baseline narrows markedly.
@@ 4.4 Ablation study
We also conduct an ablation on the output aggregation scheme: global average pooling (GAP) versus multi-head attention pooling (MAP), across four capacity levels of LBVimin table 4 right. For the tiny, 300 and small variants, GAP delivers a 0.7%, 0.4% and 0.2% higher top-1 accuracy, respectively. It is clear that when the model scales up in dimension, the advantage of GAP narrows. When the model scales to LBVim-528, the MAP gains 1.1% at a negligible 0.7 % slowdown over GAP, indicating that the expressive benefit of learned pooling emerges only when sufficient parameters are available. This size-dependent trend aligns with prior observations that simple pooling is preferable for some compact models (pan2021scalable), whereas attention-based aggregation becomes advantageous in larger models (dosovitskiy2020image_vit).
@@
4.5 Generalizability to SOTA methods
In addition to its strong performance on the Vim baseline, LBMamba demonstrates robust generalizability when applied to other SOTA models. We integrated LBMamba with four diverse architectures: VMamba (liu2024vmamba), LocalVim (huang2024localmamba), PlainMamba (yang2024plainmamba), and Adventurer (wang2025adventurer), designating the enhanced versions as LBVMamba, LBLocalVim, LBPlainMamba, and LBAdventurer, respectively. For the bidirectional models (VMamba, LocalVim, and PlainMamba), we applied both LBMamba and the sequence reversal operation. For the unidirectional Adventurer, only LBMamba was applied. Each enhanced model was configured to ensure its throughput remained comparable to its respective baseline, facilitating a fair comparison. Notably, while the original LocalVim employs a search for the optimal scanning direction ("w. search"), our LBLocalVim does not. Therefore, the performance of LBLocalVim-T should be considered a lower bound. As shown in Table 5, our method consistently improves the Top-1 accuracy by 0.5% to 3.4% across all baselines, highlighting its effectiveness as a general-purpose enhancement for Mamba-based models. Detailed experimental configurations are provided in Appendix E.
Table 5: Top-1 accuracy comparison on the ImageNet-1K dataset after applying LBMamba to four SOTA baselines. Models are evaluated at comparable throughput (T.P.) to demonstrate performance gains. Adventurer is a unidirectional scan model, while the others are bidirectional. "w./w.o. search" indicates whether the LocalVim model uses its scan direction search.
Method #Param. T.P. Top-1 acc. Method #Param. T.P. Top-1 acc.
VMamba-Nano 7M 998 78.1 PlainMamba-L1 7M 385 77.9
LBVMamba-Nano 9M 1003 79.2 LBPlainMamba-L1 15M 385 80.6
LocalVim-T (w.o. search) 8M 331 75.8 Adventurer 12M 1555 78.2
LocalVim-T (w. search) 8M 331 76.2 LBAdventurer 12M 1545 78.7
LBLocalVim-T (w.o. search) 15M 356 79.6
4.6 WSI classification
To verify LBMamba beyond natural images, we embed it into the SOTA Multiple-Instance Learning (MIL) framework, MambaMIL and SRMambaMIL (yang2024mambamil), and name them LBMambaMIL and SRLBMambaMIL, respectively. We evaluate them on 3 public available Whole Slide Image datasets, PANDA (prostate grade assessment) (bulten2022artificial_panda), TCGA-NSCLC (adenocarcinoma vs. squamous lung cancer) and TCGA-BRCA (breast invasive carcinoma sub-typing) (tcga). Dataset and training details are listed in Appendix E and F. As shown in table 6, with an additional locally backward scan, LBMambaMIL generally performans better than MambaMIL, achieves up to 3.06% higher accuracy, up to 3.39% higher F1 and up to 1.67% higher AUC. Similar improvement is also observed on SRMamba, SRLBMambaMIL achieves achieves up to 2.98% higher accuracy, up to 2.90% higher F1 and up to 1.48% higher AUC, demonstrating that the LBMamba is able to improve the performance of unidirectional scans. It also shows that our method generalize well on Giga-pixel pathology images.
@@ 4.6 WSI classification
AccAcc $F1$ AUCAUC AccAcc $F1$ AUCAUC AccAcc $F1$ AUCAUC
@@
The Effective Receptive Field (ERF) (luo2016understanding_erf) refers to the region in the input space that contributes to the activation of a specific output unit. We conduct a comparative analysis of the central pixel’s ERF on DeiT, Vim and LBVim at tiny and small scale, both before and after training. The results presented in figure 5 illustrate that DeiT shows global ERFs but it suffers from the quadratic complexity of self attention. Vim shows global ERFs and LBVim also shows global ERFs, proving that LBMamba with sequence reversing operation does not have a side effect on ERF.
Figure 5: Comparison of Effective Receptive Fields (ERF) (luo2016understanding_erf) on DeiT, Vim and LBVim at tiny and small scale. Pixels with higher intensity indicate larger responses related to the central pixel.
@@ 5 Conclusion
Limitation. LBMamba is less effective when a dedicated class token is appended to the sequence. We systematically evaluated four common variants: head class token (prepended), middle class token (inserted at the sequence midpoint), and double class token (both prepended and appended) strategy, but none of them achieves better performance than a simple global average pooling (see Appendix C). A plausible explanation is that the local backward scan treats the class token as an ordinary feature vector whose receptive field is confined to its local window; this reinforces local patterns while diluting the holistic summary that the token is meant to capture. This also limits the application of LBMamba in tokens similar to the class token, like the registers in Mamba-R (wang2025mamba_mamba_r). Unifying local bi-directionality with an effective global summarization mechanism therefore remains an open research question, and we leave a deeper investigation of this to future work.
Acknowledgement
This work was partially supported by USA NSF grants IIS-2123920 [D.S], IIS-2212046 [D.S], IIS-1715985 [H.Q], IIS-1812606 [H.Q], NSERC-DG RGPIN‐2022‐05378 [M.S.H] and Amazon Research Award [M.S.H].
@@ Appendix A LBVim block algorithm
1:  /* normalize the input sequence $\mathbf{T}_{l-1}^{\prime}$ */
2:  $\mathbf{T}_{l-1}^{\prime}$ : $(\mathtt{B},\mathtt{L},\mathtt{D})$ $\leftarrow$ $\mathbf{Norm}(\mathbf{T}_{l-1})$
3:  $\mathbf{x}$ : $(\mathtt{B},\mathtt{L},\mathtt{E})$ $\leftarrow$ $\mathbf{Linear}^{\mathbf{x}}(\mathbf{T}_{l-1}^{\prime})$
4:  $\mathbf{z}$ : $(\mathtt{B},\mathtt{L},\mathtt{E})$ $\leftarrow$ $\mathbf{Linear}^{\mathbf{z}}(\mathbf{T}_{l-1}^{\prime})$
@@
6:  $\mathbf{x}^{\prime}$ : $(\mathtt{B},\mathtt{L},\mathtt{E})$ $\leftarrow$ $\mathbf{SiLU}(\mathbf{Conv1d}_{o}(\mathbf{x}))$
7:  $\mathbf{B}$ : $(\mathtt{B},\mathtt{L},\mathtt{N})$ $\leftarrow$ $\mathbf{Linear}^{\mathbf{B}}(\mathbf{x}^{\prime}_{o})$
8:  $\mathbf{C}$ : $(\mathtt{B},\mathtt{L},\mathtt{N})$ $\leftarrow$ $\mathbf{Linear}^{\mathbf{C}}_{o}(\mathbf{x}^{\prime})$
9:  /* softplus ensures positive $\mathbf{\Delta}_{o}$ */
10:  $\mathbf{\Delta}$ : $(\mathtt{B},\mathtt{L},\mathtt{E})$ $\leftarrow$ $\log(1+\exp(\mathbf{Linear}^{\mathbf{\Delta}}(\mathbf{x}^{\prime})+\mathbf{Parameter}^{\mathbf{\Delta}}))$
11:  /* shape of $\mathbf{Parameter}^{\mathbf{A}}$ is $(\mathtt{E},\mathtt{N})$ */
12:  $\overline{\mathbf{A}}$ : $(\mathtt{B},\mathtt{L},\mathtt{E},\mathtt{N})$ $\leftarrow$ $\mathbf{\Delta}\bigotimes\mathbf{Parameter}^{\mathbf{A}}$
13:  $\overline{\mathbf{B}}$ : $(\mathtt{B},\mathtt{L},\mathtt{E},\mathtt{N})$ $\leftarrow$ $\mathbf{\Delta}\bigotimes\mathbf{B}$
14:  /* initialization with 0 */
15:  $h^{f}$ : $(\mathtt{B},\mathtt{E},\mathtt{N})$ $\leftarrow$ zeros $(\mathtt{B},\mathtt{E},\mathtt{N})$
16:  $h^{b}$ : $(\mathtt{B},\mathtt{E},\mathtt{N})$ $\leftarrow$ zeros $(\mathtt{B},\mathtt{E},\mathtt{N})$
17:  $\mathbf{y}$ : $(\mathtt{B},\mathtt{L},\mathtt{E})$ $\leftarrow$ zeros $(\mathtt{B},\mathtt{L},\mathtt{E})$
@@
32:  /* get gated $\mathbf{y}$ */
33:  $\mathbf{y}^{\prime}$ : $(\mathtt{B},\mathtt{L},\mathtt{E})$ $\leftarrow$ $\mathbf{y}\bigodot\mathbf{SiLU}(\mathbf{z})$
@@
35:  $\mathbf{T}_{l}$ : $(\mathtt{B},\mathtt{L},\mathtt{D})$ $\leftarrow$ $\mathbf{Linear}^{\mathbf{T}}(\mathbf{y}^{\prime})+\mathbf{T}_{l-1}$
36:  $\mathbf{T}_{l}^{rev}$ : $(\mathtt{B},\mathtt{L},\mathtt{D})$ $\leftarrow$ $\mathbf{Reverse}(\mathbf{T}_{l})$
@@ Appendix C Ablation on the class token
Table 8 shows several common strategies (vim) for incorporating a class token into LBVim-Ti on the ImageNet-100 dataset. Consistent with our observations on ImageNet-1K (section 4.1), LBVim-Ti under performs Vim-Ti by 0.42% as it can not process the class token. Introducing a head class token, which is prepended to the patch sequence (ViT style (dosovitskiy2020image_vit)), reduces accuracy by 3.22%. Adding a second tail token (double class token) offers a recovery (0.58%) but remains 2.64% lower than GAP. Placing the token in the middle (the strategy used by vim) of the sequence partially alleviates the degradation, yet it still under performs the no-token design. We believe this is because the class token is treated as an ordinary feature vector, its ability to aggregate global context is diluted, leading to weaker final representations. In contrast, GAP aggregates information from all positions without introducing additional parameters or disrupting the scan pattern, making it a more compatible summarization mechanism for LBMamba-based models.
@@
Appendix D Implementation details of experiments on natural images classifications
The VMmaba-Nano is the same the VMamba-tiny except a 48 initial dimensionality and the number of layer for each block is 2, 2, 6, 2. LBVMamba-Nano only has two scans in every layer and we empirically scale it only by number of layers in the third block to 11, which might not be optimal. LBPlainMamba-L1 scales the 192 dimension of PlainMamba-L1 to 288. LBLocalVim-T uses only horizontal scan and vertical scan and scale the 192 dimension of LocalVim-T to 300. All the other configurations and hyper-parameters are kept the same as their corresponding baselines. Due to a GLIB_C issue, the throughput of Adventurer and LBAdventurer is measured on a different device.
Appendix E Implementation details of WSI experiments